# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ F.2 Subgroup Analysis
We estimate the ad effect in 20 subgroups of gender, urbanicity, and education attainment for voters within the same ZIP-code area. Results are shown in Figure F.4. When $\gamma_{1}=\gamma_{0}=0$, i.e., under transportability, the effects are in general higher for non-female voters than female voters, and higher for urban voters than rural voters. When $\gamma_{1}\gamma_{0}\neq 0$, after calibration, none of the subgroups are sensitive for a significant ad effect.
Figure F.4: Heterogeneous treatment effects by the interaction between gender, urbanicity, and education subgroups while including only the NC and AZ voters in the source data. The solid lines and dashed lines represent 95% CIs under transportability for the OR estimator and the EIF estimator, respectively. The gray squares represent that all subgroups are insensitive for a significant effect.
@@ Appendix G Simulations
After generating datasets, the propensity score $\pi({\bf x})$ is estimated with the average proportion of treated units within each. The outcome regression functions $\mu_{a}({\bf x})$ and $\rho_{a}({\bf v})$ are estimated by reweighing samples with $S_{i}=1$ and $A_{i}=a$ as in (4.2). The density ratio $w({\bf v})$ is estimated with (C.1). For the OR estimator, the inference is based on 1000 bootstrap iterations. For the EIF-based estimator, the inference is based on the cross-fitting procedure with $K=2$ splits. The confidence level is set to $0.951-\alpha=$ 0.95. Simulation results are based on 1000 replicates.
@@
$\gamma_{1}=0$ Scenario (A) Scenario (B)
@@
$\gamma_{1}=0.05$ Scenario (A) Scenario (B)

### v2

@@ F.2 Subgroup Analysis
We estimate the ad effect in 20 subgroups of gender, urbanicity, and education attainment for voters within the same ZIP-code area. Results are shown in Figure F.3. When $\Gamma_{1}=\Gamma_{0}=1$, i.e., under transportability, the effects are in general higher for non-female voters than female voters, and higher for urban voters than rural voters. When $\Gamma_{0}\neq 1$ or $\Gamma_{1}\neq 1$, after calibration, none of the subgroups are sensitive for a significant ad effect.
Figure F.3: Subgroup analysis by the interaction between gender, urbanicity, and education subgroups while including only the NC and AZ voters in the source data. The solid lines and dashed lines represent 95% CIs under transportability for the OR estimator and the EIF-based estimator, respectively. The gray squares represent that all subgroups are insensitive for a significant effect.
@@ Appendix G Simulations
After generating datasets, the propensity score $\pi({\bf x})$ is estimated with the average proportion of treated units within each. The outcome regression functions $\mu_{a}({\bf x})$ and $\rho_{a}({\bf v})$ are estimated by reweighing samples with $S_{i}=1$ and $A_{i}=a$ as in (5). The density ratio $w({\bf v})$ is estimated with (C.2). For the OR estimator, the inference is based on 1000 bootstrap iterations. For the EIF-based estimator, the inference is based on the cross-fitting procedure with $K=2$ splits. The confidence level is set to $0.951-\alpha=$ 0.95. Simulation results are based on 1000 replicates.
@@
$\gamma_{1}=1$ Scenario (A) Scenario (B)
@@
$\gamma_{1}=1.05$ Scenario (A) Scenario (B)